The width of a rectangular painting is tw thirds of its length. If the perimeter is 60 inches then what are the length and width?
Answer by checkley77(12844)

W=2/3L
2W+2L=60
2(2L/3)+2L=60
4L/3+2L=60
(4L+3*2L)/3=60
(4L+6L)/3=60
10L/3=60
10L=3*60
10L=180
L=180/10
L=18 ANS.
W=2/3*18
W=36/3=12 ANS.
PROOF:
2*12+2*18=60
24+36=60
60=60